• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 20:00:29 on July20,2005

    Entry number 148756

    keyword=spectrometer runs

    Spectrometer run 2176 to see where we are.

    HV on S0
    L-arm: -1400 and -1550
    R-arm: Both at -1550

    HAPPEX Detector HV
    L-arm -1900
    R-arm -1800

    Beam current 2 uA
    T1 = 6x10^4 Hz
    T2 = 6000
    T3 = 5x10^4 Hz
    T4 = 4300
    (Note, no huge difference L-R for T1 vs T3, or T2 vs T4 !!)
    T5 = T6 = 0
    T7 = 9.6 x 10^4 Hz
    T8 = 5.3 x 10^4 Hz
    (More of a difference T7,T8 but...)

    Run 2177: New T0 for Triggers T2 and T4. Decreased their
    delay by 32 nsec to reduce the gap to the S0 trigger.

    Checked that T0 shifted the right way.
    There is still a gap of 50 nsec between the triggers, which is
    pretty big, but acceptable. See figs 1 and 2.

    Next step:
    We ask for beam to move to steer more through the middle of
    the spectrometer system.
    Ask for BPM4A at (x,y) = (1.6,0)
    and BPM4B at (x,y) = (-1,4,0)
    MCC is trying to achieve this now....

    FIGURE 1

    FIGURE 2